OpenStudy (anonymous):

A force of 20N gives an object an acceleration of 5 m/s^2. a. What force would be needed to give the same object an acceleration of 1.0 m/s^2 ? b. What force would be needed to give an acceleration of 10 m/s^2 ?

OpenStudy (anonymous):

are you familiar with Newton's second law?

OpenStudy (anonymous):

Yes, I'm just not sure how to solve this.

OpenStudy (anonymous):

Newton's second law states that force = rate of change of momentum. and upon simplification yields the result Force = mass * acceleration. since you are given the force and the acceleration solve for the mass (m) F= m*a 20 = m* 5 and then use this mass and the same equation to find the acceleration in part a) of the question.

OpenStudy (anonymous):

part a) Force = m * 1 = _________Newtons part b ) Force = m* 10 = __________ Newtons

OpenStudy (anonymous):

did you understand the method?

OpenStudy (anonymous):

yes. thank you!

OpenStudy (anonymous):

Its been a pleasure :)